# Entropy-Lens: The Information Signature of Transformer Computations

### v2

@@ B.1 Entropy-Lens for Vision Transformers
We hypothesize a phase transition in entropy behavior as model size increases, occurring somewhere between 8787M and 307307M parameters.
@@
Appendix C Evaluation Details
@@
In all cases, prompt design plays a critical role in ensuring robust comparisons across experimental conditions. To this end, we employed multiple prompt variations. The subsections below report the full set of templates used, as referenced in Sections 5.2, 5.3, and 5.5 of the main paper.
@@
C.1 Prompt templates for TinyStories tasks
In Section 5.2, we describe an experimental setup designed to test whether entropy profiles can identify different types of tasks. To this end, we used the TinyStories dataset (Eldan and Li 2023) and constructed prompts combining short stories with specific task instructions. Each task type—generative, syntactic, and semantic—was associated with two distinct natural language formulations, referred to as task prompts. These are listed in Table 4. By varying the task prompt, we ensure that our classification results are not simply driven by surface-level textual artifacts, but instead reflect deeper computational signatures captured by the entropy profiles.
This table complements the prompt templates (base, reversed, and scrambled) described in the main text and provides the full set of instructions used to elicit different model behaviors.
@@ C.1 Prompt templates for TinyStories tasks
Table 4: Prompt templates used for TinyStories tasks.
C.2 Prompt templates for MMLU correctness classification
In Section 5.3, we test whether entropy profiles can distinguish between correct and incorrect answers produced by language models. To construct the dataset for this experiment, we used the Massive Multitask Language Understanding (MMLU) benchmark (Hendrycks et al. 2021), applying three distinct prompt styles to elicit different answer behaviors from the models. Table 5 reports the full prompt templates used in this experiment. Each template presents the same multiple-choice question in a different instructional format: the Base format presents the question directly; the Instruct format introduces an explicit instruction to select a single correct answer; and the Humble format includes an additional fallback directive encouraging the model to guess randomly if uncertain.
This variation in prompting allows us to control for instruction framing and to evaluate whether entropy profiles can capture response confidence and correctness robustly across different model behaviors. The table shown here complements the description in the main text.
@@ C.2 Prompt templates for MMLU correctness classification
Table 5: Prompt templates used for the MMLU dataset.
C.3 Prompt Construction for the Topic-Format Dataset
@@
C.4 Experimental and hardware setup
All experiments were conducted on a compute node equipped with an NVIDIA L40 GPU, an Intel Xeon Gold CPU, 128 GB of RAM, and running Ubuntu 22.04. The primary software frameworks used were PyTorch, Transformer-Lens, and HuggingFace Transformers. Inference on LLMs was performed using float16 precision for improved efficiency.
@@ D.1 Minimal Entropy Profile Extraction
The code in listing 1 demonstrates how to compute an entropy profile for a single prompt using a standard decoder-only transformer. It relies only on model forward passes and the use of logit-lens-style decoding. No gradients, fine-tuning, or access to internal model weights are required.

### v3

@@ A.4 Entropy-Lens for Vision Transformers
We hypothesize a phase transition in entropy behavior as model size increases, occurring somewhere between 8787M and 307307M parameters, though a more extensive study would be required to confirm this hypothesis.
@@
A.5 Stability across Rényi entropy variants
To qualitatively explore how the Rényi entropy affects the structure of entropy profiles, we compute pairwise cosine similarities between profiles generated with different values of $\alpha$. This analysis is performed on a subset of the topic-format dataset (see Appendix B.4), and the resulting similarity matrices are shown in Figure 9.
We observe three distinct regimes as $\alpha$ varies: (1) For very small values of $\alpha$ (e.g., $\alpha<0.2$), the similarity matrices are nearly flat, with profiles being almost identical across examples. This is expected, as Rényi entropy in this regime weights all tokens with non-zero probability almost equally, and usually all tokens have non-null probability. (2) For large values of $\alpha$ (e.g., $\alpha>20$), the similarity matrices also flatten. In this case, the entropy becomes increasingly dominated by the few tokens with highest probabilities. Since these sets of top tokens tend to have similar cardinalities (in the limit equal to 1), the profiles collapse into a narrow set of values, losing expressiveness and becoming more sensitive to local fluctuations. (3) Between these extrema lies an informative regime, approximately $0.5\leq\alpha\leq 20$, where entropy profiles are heterogeneous enough to retain meaningful variation. This is reflected in the standard deviation of the similarity matrices, which peaks in this interval.
We find that the qualitative structure of entropy profiles, shown in Figure 9, and the diagnostic results reported in Section B.4 are stable when replacing Shannon entropy with Rényi entropies for a broad range of $\alpha$ values.
Appendix B Evaluation Details
@@
In all cases, prompt design plays a critical role in ensuring robust comparisons across experimental conditions. To this end, we employed multiple prompt variations. The subsections below report the full set of templates used, as referenced in Sections 5.2, A.2, and 5.3 of the main paper.
@@
B.1 Model family classification
In Section 5.1, we show how entropy profiles can effectively identify both model families and model sizes. Our analysis reveals that entropy profiles exhibit qualitatively distinct patterns across different model families and sizes, as illustrated through scatterplot visualizations. By applying t-SNE dimensionality reduction, we cluster models by family, indicating that entropy profiles capture meaningful structural differences between architectures. To quantitatively assess the classification capabilities of entropy profiles, we employ a k-nearest neighbors classifier ($k=3$ and euclidean distance) to predict both model families and sizes based on their entropy traces. The classification results are presented in Table 6. To obtain labeled model size categories, we binned models into 4 classes based on parameter count (in billions): <<1B, 1-3B, 3-5B, and >>5B.
Table 6: F1-scores for model family and model size classification. Each reported value is the mean across 10 runs, with the standard deviation computed over random 50/50 train–test splits.
Task Macro F1-score
[…]
Preprocessing Steps.
Since entropy traces vary in length across models due to different layer counts, we apply linear interpolation to standardize all traces to the same length. Additionally, to ensure fair classification performance, we standardize the samples to reduce bias from scaling effects in the entropy profiles, allowing the classifier to focus on the characteristics of each trace.
B.2 Prompt templates for TinyStories tasks
In Section 5.2, we describe an experimental setup designed to test whether entropy profiles can identify different types of tasks. To this end, we use the TinyStories dataset (Eldan and Li, 2023) and construct prompts combining short stories with specific task instructions. Each task type, generative, syntactic, and semantic, was associated with two distinct natural language formulations, referred to as task prompts. These are listed in Table 7. By varying the task prompt, we ensure that our classification results are not simply driven by surface-level textual artifacts, but instead reflect deeper computational signatures captured by the entropy profiles. The same table complements the prompt templates (base, reversed, and scrambled) described in the main text and provides the full set of instructions used to elicit different model behaviors. Figure 10 shows the mean and standard deviation of the entropy profiles for different tasks.
Figure 10: Average entropy profiles with shaded standard deviation for different task types: generative, syntactic, and semantic. These tasks are induced with the prompts described in Appendix B.2. Left: Llama-3.2-it. Right: Gemma-2-it.
Table 7: Prompt templates used for TinyStories tasks.
@@ B.2 Prompt templates for TinyStories tasks
B.3 Prompt templates for MMLU correctness classification
In Section A.2, we test whether entropy profiles can distinguish between correct and incorrect answers produced by language models. To construct the dataset for this experiment, we used the Massive Multitask Language Understanding (MMLU) benchmark (Hendrycks et al., 2021), applying three distinct prompt styles to elicit different answer behaviors from the models. Table 8 reports the full prompt templates used in this experiment. Each template presents the same multiple-choice question in a different instructional format: the Base format presents the question directly; the Instruct format introduces an explicit instruction to select a single correct answer; and the Humble format includes an additional fallback directive encouraging the model to guess randomly if uncertain.
This variation in prompting allows us to control for instruction framing and to evaluate whether entropy profiles can capture response confidence and correctness robustly across different model behaviors. Table 8 complements the description in the main text.
Table 8: Prompt templates used for the MMLU dataset.
@@ B.3 Prompt templates for MMLU correctness classification
B.4 Prompt Construction for the Topic-Format Dataset
@@
B.5 Experimental and hardware setup
The experiments were conducted on:
  • A compute node equipped with an NVIDIA L40 GPU, an Intel Xeon Gold CPU, 128 GB of RAM, and running Ubuntu 22.04.
  • A compute node equipped with an NVIDIA H100 GPU, an Intel Xeon Platinum CPU, 240 GB of RAM and running Ubuntu 22.04.
The primary software frameworks used were PyTorch, Transformer-Lens, and HuggingFace Transformers. Inference on LLMs was performed using float16 precision for improved efficiency. Part of the experiments were conducted on Chameleon Cloud (Keahey et al., 2020). Source code is available at: github.com/christopher-irw/Entropy-Lens.
Appendix C Theoretical Considerations
In addition to the empirical results presented in the main text, we provide here a preliminary theoretical perspective that connects entropy profiles to the literature on memorization in transformers. Our goal is not to give a full formal treatment, but rather to outline how previous definitions of memorization, based on information theory, can be related to the quantities we compute. We first recall the frameworks introduced by Brown et al. (2021) and Morris et al. (2025), and then show how our entropy profiles can be interpreted as layer-wise estimates of memorization.
C.1 Entropy and Memorization
From Shannon to Kolmogorov memorization.
Brown et al. (2021) introduced an information-theoretic framework to quantify memorization in trained models. Given a training data distribution $X$, a family of data-generating processes $\Theta$, and a training algorithm $L:X\mapsto\hat{\Theta}$ mapping training sets to trained models, they define memorization as the mutual information between $X$ and $\hat{\Theta}$
[…]
This quantity captures how much information about $X$ is retained in the distribution over trained models. It can be decomposed into
$\text{mem}(X,\hat{\Theta})=\text{mem}_{I}(X,\hat{\Theta},\Theta)+\text{mem}_{U}(X,\hat{\Theta},\Theta),$ (5)
where $\text{mem}_{I}$ measures generalization and $\text{mem}_{U}$ the unintended memorization (i.e., information about $X$ not attributable to the process $\Theta$).
Building on this formalism, Morris et al. (2025) extended the analysis from distributions to individual instances, moving from Shannon to Kolmogorov information theory. The Kolmogorov complexity of an instance $x$ given model parameters $\hat{\theta}$ is
$H^{k}(x|\hat{\theta})=\min_{s}\{|s|:f(s,\hat{\theta})=x\},$ (6)
where $f$ is a computational model (imagine a decoder) that can take as input $x$ and $\theta$. The exact definition of Kolmogorov complexity is not computable in general. In practice, it can be approximated via arithmetic coding as
[…]
Connecting to entropy profiles.
[…]
Now, rather than computing this value only for the full model, we do so for every intermediate truncation. Let $\hat{\Theta}_{i}$ denote the sub-model obtained by retaining only the first $i$ layers of the trained transformer and applying an early exit. The corresponding conditional entropies are
$H(X|\hat{\Theta}_{i}),\quad i=1,\dots,N,$ (8)
and the sequence $\{H(X|\hat{\Theta}_{i})\}_{i=1}^{N}$ constitutes the entropy profile. This is equivalent to
$H(X|\hat{\Theta}_{i})=H(X)-I(X,\hat{\Theta}_{i}),$ (9)
i.e. the negative mutual information between the dataset and the truncated model up to a constant $H(X)$, which is equal for all layers.
This perspective suggests that entropy profiles can be interpreted as measuring how memorization is distributed across depth. Crucially, our empirical results show that this allocation does not follow a simple monotonic trend, as one might have expected a priori. Instead, it varies in a systematic way depending on model family, task, format, and confidence, revealing non-trivial patterns of information storage that had not been documented before.
@@ D.1 Minimal Entropy Profile Extraction
The code in Listing LABEL:lst:entropy-signature demonstrates how to compute an entropy profile for a single prompt using a standard decoder-only transformer. It relies only on model forward passes and the use of logit-lens-style decoding. No gradients or fine-tuning are required, but only forward access to intermediate hidden states and the output head.